# CompoDiff: Versatile Composed Image Retrieval With Latent Diffusion

### v3

@@ Appendix B Full experiment results of Table 3
${}^{\dagger}$ 30.70 50.43 33.52 46.54 35.49 47.01 33.24 47.99 40.62
${}^{\dagger}$ 32.32 51.65 34.92 48.38 35.65 48.10 34.30 49.38 41.84
CompoDiff (${}^{\dagger}$ 37.69 49.08 32.24 46.27 38.12 50.57 36.02 48.64 42.33
CompoDiff ${}^{\ddagger}$ (${}^{\dagger}$ 38.10 52.48 33.91 47.85 40.07 52.22 37.36 50.85 44.11
CompoDiff (${}^{\dagger}$ 41.31 55.17 37.78 49.10 44.26 56.41 39.02 51.71 46.85
@@
${}^{\dagger}$ 32.17 53.32 34.80 48.10 36.58 47.63 34.52 49.68 42.10
${}^{\dagger}$ 37.21 60.71 42.75 60.50 42.98 65.49 40.98 62.23 51.61
CompoDiff (${}^{\dagger}$ 40.88 53.06 35.53 49.56 41.15 54.12 39.05 52.34 46.31
CompoDiff (ViT-L + ${}^{\dagger}$ 40.65 57.14 36.87 57.39 43.93 61.17 40.48 58.57 49.53
CompoDiff (${}^{\dagger}$ 41.68 56.02 38.39 51.03 45.70 57.32 39.81 51.90 47.73
@@
${}^{\dagger}$ 12.75 33.84 47.75 80.20 21.95 43.88 62.06 17.35
${}^{\dagger}$ 12.82 36.83 48.19 81.91 24.12 46.47 63.07 18.47
CompoDiff (${}^{\dagger}$ 18.24 53.14 70.82 90.25 57.42 77.10 87.90 37.83
CompoDiff ${}^{\ddagger}$ (${}^{\dagger}$ 19.37 53.81 72.02 90.85 59.13 78.81 89.33 39.25
CompoDiff (${}^{\dagger}$ 26.71 55.14 74.52 92.01 64.54 82.39 91.81 45.63
@@
${}^{\dagger}$ 18.85 51.44 68.01 91.93 38.85 62.00 77.68 28.85
${}^{\dagger}$ 39.99 73.63 86.77 96.55 68.41 86.12 94.80 54.20
CompoDiff (${}^{\dagger}$ 21.30 55.01 72.62 91.49 58.82 77.60 88.37 40.06
CompoDiff (ViT-L + ${}^{\dagger}$ 22.35 54.36 73.41 91.77 62.55 81.44 90.21 42.45
CompoDiff (${}^{\dagger}$ 32.39 57.61 77.25 94.61 67.88 85.29 94.07 50.14

### v4

@@ Appendix B Full experiment results of Table 3
ARTEMIS 30.70 50.43 33.52 46.54 35.49 47.01 33.24 47.99 40.62
CLIP4Cir 32.32 51.65 34.92 48.38 35.65 48.10 34.30 49.38 41.84
CompoDiff (ViT-L) 37.69 49.08 32.24 46.27 38.12 50.57 36.02 48.64 42.33
CompoDiff (ViT-L) 38.10 52.48 33.91 47.85 40.07 52.22 37.36 50.85 44.11
CompoDiff (ViT-G) 41.31 55.17 37.78 49.10 44.26 56.41 39.02 51.71 46.85
@@
ARTEMIS 32.17 53.32 34.80 48.10 36.58 47.63 34.52 49.68 42.10
Combiner 37.21 60.71 42.75 60.50 42.98 65.49 40.98 62.23 51.61
CompoDiff (ViT-L) 40.88 53.06 35.53 49.56 41.15 54.12 39.05 52.34 46.31
CompoDiff (ViT-L + T5-XL) 40.65 57.14 36.87 57.39 43.93 61.17 40.48 58.57 49.53
CompoDiff (ViT-G) 41.68 56.02 38.39 51.03 45.70 57.32 39.81 51.90 47.73
@@
ARTEMIS 12.75 33.84 47.75 80.20 21.95 43.88 62.06 17.35
CLIP4Cir 12.82 36.83 48.19 81.91 24.12 46.47 63.07 18.47
CompoDiff (ViT-L) 18.24 53.14 70.82 90.25 57.42 77.10 87.90 37.83
CompoDiff (ViT-L) 19.37 53.81 72.02 90.85 59.13 78.81 89.33 39.25
CompoDiff (ViT-G) 26.71 55.14 74.52 92.01 64.54 82.39 91.81 45.63
@@
ARTEMIS 18.85 51.44 68.01 91.93 38.85 62.00 77.68 28.85
Combiner 39.99 73.63 86.77 96.55 68.41 86.12 94.80 54.20
CompoDiff (ViT-L) 21.30 55.01 72.62 91.49 58.82 77.60 88.37 40.06
CompoDiff (ViT-L + T5-XL) 22.35 54.36 73.41 91.77 62.55 81.44 90.21 42.45
CompoDiff (ViT-G) 32.39 57.61 77.25 94.61 67.88 85.29 94.07 50.14